OpenStudy (anonymous):

The speed of a passanger train is 6 mph faster than a freight train. The passanger train travels at 290 miles in the same time it takes the freight train to travel 260 miles. Find the speed of each train. Passanger? Freight?

OpenStudy (anonymous):

speed of Passenger Train =P Frieght Train=F F+6=P

OpenStudy (anonymous):

P*t=290 f*t=260

OpenStudy (anonymous):

good from here?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

P*t=290 f*t=260 so we have to find p , and f

OpenStudy (anonymous):

we can rewrite p in terms of f p=f+5 (f+5)t=290 ft+5t=290

OpenStudy (anonymous):

we know what ft is ft=260 5t=290-ft 5t=30 t=6

OpenStudy (anonymous):

Let p be the speed of the passenger train. Solve the following simultaneous equations for p and t:\[\{p*t =290,(p-6)*t=260\} \]\[\{p\to 58,t\to 5\} \]Passenger Train speed = 58 mph Freight Train speed = 52 mph